### 28 Apr 2026 v2

# SketchVLM: Vision language models can annotate images to explain thoughts and guide users

Abstract
[…]
**footnotetext: All three co-first authors made major contributions to dataset creation, experiments, and manuscript. See Sec.˜7 for a detailed author contribution statement.
1 Introduction
From browsers (Atlas, MS Edge [39]) to office products (MS Copilot, Gemini [43]), the market for LLM-powered chatbots for general end users is estimated to grow from \$1.1B in 2023 to \$83B in 2032 [6]. As these assistants increasingly answer image-based questions in high-frequency consumer and productivity workflows [39, 35, 43, 6], users need responses they can quickly understand and verify [29].
However, modern VLMs such as Gemini-3-Pro and GPT-5 typically respond with a block of text, which can be difficult to verify [55] (Fig.˜A2). To explain their thoughts on an image problem, humans often point, circle, underline, label, and annotate directly on the image. For example, when a user asks to check a car’s oil level, a visual annotation is easier to verify than a paragraph of text (Fig.˜1) [26, 31].
Some VLMs emit point coordinates for referencing objects (e.g., MoonDream [45], Molmo [2]), but they do not support free-form visual annotation. Other models are trained to generate visual reasoning traces but do not generalize well beyond their training domains [48, 18]. Image-editing models can visualize intermediate reasoning steps for multimodal questions [57], but risk altering the source image in unintended ways, undermining user trust [41].
Figure 1: For complex questions, modern chatbots like ChatGPT often return long text responses (a) that are hard for users to understand, verify, and follow. In contrast, SketchVLM guides users (b) step-by-step by annotating the input image and grounding answers to relevant image regions—here, guiding a user how to check their car’s oil level (source: https://www.youtube.com/watch?v=tNNyu9S65E4).
To address these issues, we propose SketchVLM, a state-of-the-art (SotA) system that draws SVG annotations in a separate layer overlaid on top of the input image to explain its reasoning. SketchVLM grounds its annotations directly on the original image without modifying it and without requiring task-specific training. We test SketchVLM with multiple VLM backbones, including Gemini-3-Pro-Preview ([Uncaptioned image]) [36] and GPT-5 ([Uncaptioned image]) [30]. We collectively refer to these models harnessed with SketchVLM as SketchVLMs and individually as [Uncaptioned image] sketchVLM and [Uncaptioned image] sketchVLM. We compare against the leading alternative approaches for generating visual annotations such as the SotA image-editing model Nano Banana Pro ([Uncaptioned image]) [13] and specialized VLMs fine-tuned to produce image annotations, ViLaSR ([Uncaptioned image]) [48] and ThinkMorph ([Uncaptioned image]) [18].
On a comprehensive evaluation of seven tasks: (a) three drawing tasks (connecting-the-dots, labeling parts of an object, and drawing shapes around objects) and (b) four visual reasoning tasks (two physics understanding tasks, one counting task and one maze navigation task), our main findings are:
Input Refer to caption sketchVLM $\bm{+} +\bm{+} Refer to caption$ Refer to caption ViLaSR [48]
(a)   Ball Drop Refer to caption Refer to caption Refer to caption Refer to caption
Which bucket will the ball fall into? Answer: 2 Refer to caption Answer: 2 Refer to caption Ball goes through platform Refer to caption Answer: 3 Refer to caption Ball goes through platform Refer to caption
Input Refer to caption sketchVLM Refer to caption ThinkMorph [18] Refer to caption ViLaSR [48]
(b)   Connect-the-Dots Refer to caption Refer to caption Refer to caption Refer to caption
Connect the dots in order. Correct Refer to caption Improper Lines Refer to caption Improper Lines Refer to caption
Input Refer to caption sketchVLM $\bm{+} +\bm{+} Refer to caption$ Refer to caption ThinkMorph [18]
(c)   Maze Navigation Refer to caption Refer to caption Refer to caption Refer to caption
Is the path “right, right, up, up” from green valid? Answer: Yes Refer to caption Answer: No Refer to caption   Wrong path Refer to caption Answer: Yes Refer to caption   Wrong path Refer to caption
Figure 2: Our Refer to caption sketchVLM (Gemini-3-Pro-Preview) draws more accurate predicted trajectories in Ball Drop (a); connects the dots more accurately (b); and sketches more plausible maze navigation paths (c). Nano Banana ($\bm{+} +\bm{+} Refer to caption$) often undesirably alters the image and draws implausible trajectories in ball drop and maze navigation. Specialist VLMs (Refer to caption and Refer to caption) fine-tuned to sketch often fail to generalize to new tasks.
  • SketchVLMs based on frontier models ([Uncaptioned image] and [Uncaptioned image]) generate annotations of superior generalizability, accuracy, and annotation quality compared to those by specialized fine-tuned sketching models ([Uncaptioned image], [Uncaptioned image]) (Secs.˜5.2, 5.3, 5.6, 5.7, 5.8, 5.9 and 5.10).
  • Nano Banana ([Uncaptioned image]) is unable to generate a separate overlay layer and frequently alters the original image when generating in-image annotations (Secs.˜5.4 and 5.5).
  • SketchVLMs have similar accuracy in single-turn and multi-turn settings, but are significantly faster with single-turn generation (Sec.˜5.11).
  • Adding an external grid of xxyy coordinates to the input image improves [Uncaptioned image] sketchVLM’s drawing capability and question answering accuracy but is not necessary for [Uncaptioned image] sketchVLM (Sec.˜5.1).
2 Related Work
Native image-editing models   (e.g., GPT-Image-1.5 and [Uncaptioned image]) can directly modify images to add annotations, but their performance can be inconsistent for reasoning-heavy tasks [57]. Open-source native multimodal autoregressive models (e.g., Chameleon [42] and Bagel [15]) enable interleaved text–image generation, but they lack an editable annotation layer aligned to the input. In contrast, SketchVLMs generate a non-destructive SVG overlay on the input image (Tab.˜1).
Tool-calling and code generation   Agentic systems improve visual reasoning by writing code that invokes external tools or manipulates the input image. V* [49] uses LLM-guided visual search to localize target objects in high-resolution images, then crops relevant regions for closer inspection. Visual Sketchpad [19] and OpenThinkIMG [40] equip VLMs with modular vision tools such as segmentation, object detection, and OCR to support multi-step reasoning. PyVision [54] generates Python code to draw structured overlays on input images across multiple turns. Other training-free methods leverage internal attention or gradient maps to automatically crop and zoom into salient regions [53]. These systems excel at fine-grained visual understanding, but they typically rely on external tools or code execution rather than prompting a single VLM to produce user-facing free-form annotations directly on the image.
Visual prompting   e.g., drawing coordinate points or horizontal lines directly on an input image can improve the vision capabilities of VLMs [50, 22, 20]. Similarly, SketchAgent [46] appends a coordinate grid to the edge of the input image to allow the model to reference precise xxyy positions in the image.
Visual sketching   Whiteboard-of-Thought [27] prompts an LLM to produce Matplotlib code rendered on a blank canvas to give the model a space to draw before responding to text-based questions. D2R [33] interleaves textual chain-of-thought with rendered visual drafts of its proposed actions overlaid on the input image at each reasoning step, enhancing dynamic spatial reasoning across multiple turns. VDLM [47] converts images into SVG and then into a more LLM-interpretable format to improve visual understanding. SketchAgent [46] and SketchFormer [38] focus on sketch generation as a standalone task on a blank canvas. In contrast, SketchVLMs generate non-destructive, editable SVG annotations directly on existing input images so that users can inspect the model’s reasoning without altering the source image.
Fine-tuned sketching models   MVoT [23] fine-tunes Chameleon [42] to generate interleaved text and image reasoning traces, visualizing intermediate states to support multi-step spatial reasoning. LatentSketchpad [52] and DeepSketcher [51] both move visual reasoning into learned latent or embedding spaces. LatentSketchpad is built with Gemma3 and Qwen2.5-VL-7B, while DeepSketcher uses Qwen2.5-VL-7B. ViLaSR [48] post-trains Qwen-2.5-VL-7B to sketch on the input image with an SVG overlay before responding. Similarly, ThinkMorph is fine-tuned from BAGEL-7B-MoT [15] to generate visual sketches that support its answers, though it directly edits the input image rather than overlaying SVG annotations. Unlike these approaches, we build a harness around SotA VLMs to enable them to annotate on input images by generating a layer of SVGs. Therefore, different from the literature, our approach is training-free and enjoys the generalizability of SotA VLMs to new domains.
Table 1: Comparison of sketching models and methods. Annotation type describes the visual artifact used during reasoning: Vector overlay denotes structured, non-destructive marks (e.g., strokes/boxes/text) aligned to an image or canvas, while Image edit denotes pixel-space image modification or synthesis that may change image content. Input image is marked only when a provided image is the primary visual context (vs. blank canvas or purely generative visual thoughts). Free-form drawing indicates support for arbitrary stroke-like annotations beyond a fixed mark set.
Model name Training-free Multi-turn Input image Free-form drawing Annotation type
SketchVLM (Ours) Vector overlay
VisualSketchPad [19] Vector overlay
PyVision [54] Vector overlay
SketchAgent [46] Vector overlay
D2R [33] Image edit
Whiteboard-of-Thought [27] Vector overlay
ViLaSR [48] Vector overlay
OpenThinkIMG [40] Image edit
ThinkMorph [18] Image edit
MVoT [23] Image edit
LatentSketchpad [52] Image edit
DeepSketcher [51] Image edit
3 SketchVLM
SketchVLM combines three components: visual prompting to aid spatial reference, a system prompt that elicits structured stroke outputs, and XML-to-SVG conversion that renders those strokes as an overlay on the source image (Fig.˜1).
Visual prompting   To make VLMs draw more reliably on tasks that require precision, such as Connect-the-Dots, we follow SketchAgent [46] and append a coordinate grid to the left and bottom of each image, scaled to the image resolution (Fig.˜D12).
Input prompt   To enable VLMs to generate annotations, we introduce a system prompt (Sec.˜F.2) that instructs the model to produce stroke sequences in a specific format (e.g., XML-style <s1>, <s2>, ... <sN> tags each containing a list of points) corresponding to reasoning steps. We provide instructions for drawing primitives including rectangles, arrows, text labels, straight lines, and Bézier curves. The model is then given the task prompt (e.g., “Which bucket will the ball fall into?” (Fig.˜2)) and is instructed to annotate its reasoning on the image before responding with a final answer (Fig.˜1).
SVG conversion   We parse the model’s XML output to a standardized SVG output that can be overlaid on top of the original input image. If there are exactly two points in a stroke, we overlay a straight line. Otherwise, given a stroke described by $m$ ordered samples $S_{i}=\{(x_{j},y_{j})\}_{j=1}^{m}$ and corresponding normalized timestamps $T_{i}=\{t_{j}\}_{j=1}^{m}$ with $t_{j}\in[0,1]$, we fit a smooth Bézier curve and render it as SVG. Following SketchAgent [46], a least squares solution is found for the control points in a cubic Bézier curve. See Fig.˜F5 for an example of the stroke output and the corresponding overlay. Following SketchAgent, we run all experiments using XML format. Given that VLMs produce JSON at comparable quality and JSON is more human-readable, we use JSON for our interactive demo.
4 Evaluation
4.1 7 Tasks
1. Connect-the-Dots [Uncaptioned image]   contains 100 images spanning three subsets: 21 randomly generated dot patterns, 30 connect-the-dots puzzles derived from silhouette SVGs, and 49 worksheet-style images collected from online sources. Models must locate each dot and connect them in order (Secs.˜B.1, 2 and D.3).
2. Counting Objects [Uncaptioned image]   contains 746 images drawn from CountBench [5, 34], and Pixmo-Count [14]. We include object counts from 0 to 10 and filter out unsuitable Pixmo-Count examples. Models must count the target objects and place numbered markers on each one (Secs.˜B.2, 5 and D.4).
3. Drawing Shapes around Objects [Uncaptioned image]   uses 1,000 images selected from the 5,000-image COCO validation set [25]. We choose images to balance object count and object size across classes. Models must localize objects by drawing rectangles or ovals (Secs.˜B.3, 6 and D.5).
4. Part Labeling [Uncaptioned image]   contains 985 images selected from PACO [37] and Pascal-Part [10], covering 52 object classes. We keep images with a single target object occupying at least 10% of the image area and with at least four annotated part labels, while maintaining class balance. Models must place the correct text labels at the corresponding part locations (Secs.˜B.4, 7 and D.6).
5. Maze Navigation [Uncaptioned image]   contains 200 generated $3\times 3$ grid mazes. We vary the shortest path length from 3 to 8 steps and create invalid paths by perturbing one direction in the ground-truth path. Models must trace a proposed path and determine whether it reaches the goal without crossing walls (Secs.˜B.5, 2 and D.2).
6. VPCT   is the Visual Physics Comprehension Test [8], which contains 100 hand-crafted images. Models must predict which container a dropped ball will land in (Secs.˜B.6, 2 and D1).
7. Ball Drop [Uncaptioned image]   contains 198 synthetically generated images with harder-to-guess answers and ground-truth ball trajectories produced using PHYRE [4]. We generate equal numbers of images with 1, 2, and 3 randomly placed lines, and randomize the ball’s horizontal position. Models must predict the landing container out of four choices and trace the ball trajectory (Secs.˜B.6, 10, D3 and D2).
4.2 Setup
Single-turn and multi-turn   We evaluate SketchVLMs in (1) single-turn, where the models output all annotations and their final answer in one response, and (2) multi-turn, producing one stroke per turn, to simulate iterative, real-world conversations. During each turn, the VLM receives the image with all previous annotations rendered, and the annotations’ text representations. The model then outputs its final text answer on the last turn (Fig.˜3).
Visual prompting   We evaluate the necessity of a coordinate grid for the model to reference specific points in an image. When omitting the grid, the model outputs coordinates on a normalized $\times1000$ scale (Fig.˜D14).
(a) Single-turn
(b) Multi-turn
Figure 3: Single-turn and multi-turn generation on the same VPCT sample. In (a) single-turn, SketchVLM receives the system prompt, the task prompt, and the input image, then outputs all annotations and the final answer in a single model call. In (b) multi-turn, Turn 1 uses the same inputs and outputs one annotation. For later turns, the model reuses the system prompt, the task prompt, and the previous annotations, which are provided in both the rendered image and text form. This process repeats until the model outputs its final text answer on the last turn.
4.3 Baselines
Figure 4: Four approaches for making VLMs answer visual questions and annotate images.
(a) Refer to caption outputs text only. No drawings generated.
(b) Refer to caption sketchVLM draws on the image while outputting text.
(c) Refer to caption only edits the image.
(d) $\bm{+} +\bm{+} Refer to caption$ takes the edited image from Refer to caption and gives it to Refer to caption to respond.
We compare SketchVLMs against three baselines that represent the SotA approaches for producing visual annotations on images (Fig.˜4):
  1. 1.
    Image Editing Model + VLM: [Uncaptioned image] is an image-editing model that can annotate images, but produces no text answer. To obtain a text response, we feed its edited image to [Uncaptioned image] (denoted $\bm{+} +\bm{+} [Uncaptioned image]$).
  2. 2.
    Fine-tuned Sketching Models: [Uncaptioned image] is a model fine-tuned from Qwen-2.5-VL-7B [3] to autoregressively generate SVG annotations on the input image over multiple turns, and [Uncaptioned image] is a model fine-tuned from BAGEL-7B-MoT [15] to directly edit the input image while also producing a text answer (Tabs.˜F1 and G.1).
  3. 3.
    Default VLMs: We include text-only [Uncaptioned image] and [Uncaptioned image] as simple baselines.
4.4 Metrics
A correct text answer from the model is not enough if the annotation is uninformative, and a plausible annotation can be misleading if it contradicts the text response. We therefore measure three distinct aspects of model performance:
Accuracy   serves as the primary measure of task performance and tests the effect of sketching on answer correctness.
Annotation–text alignment   measures how faithful the visual traces are to the text answer. We ask a VLM judge [9] to infer the answer from the annotations alone and report how often the judge’s inferred answer matches the model’s final text answer.
Annotation quality   allows us to distinguish models that produce informative annotations from those that output low-quality or incoherent drawings. We adopt a VLM-as-a-Judge approach using a rubric scored from 1 to 5 (Appendix˜E) that evaluates annotation plausibility and visual clarity for each task.
5 Results
Table 2: SketchVLMs produce visual reasoning traces while maintaining competitive accuracy. $\bm{+} +\bm{+} [Uncaptioned image]$ underperforms default [Uncaptioned image], and fine-tuned sketching models ([Uncaptioned image], [Uncaptioned image]) perform near random chance on visual reasoning tasks.
Model VPCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]draw shapes [Uncaptioned image]connect dots
video physics ball drop maze trace counting labeling
RMSE Order Acc%
[Uncaptioned image] sketchVLM 96.0 $\pm$ 1.4 79.7 $\pm$ 2.8 98.0 $\pm$ 1.7 94.5 60.3 58.8 55.4 5.92 99.0
[Uncaptioned image] 89.3 $\pm$ 2.2 83.8 $\pm$ 3.4 99.3 $\pm$ 0.8 93.0 64.1 63.1 59.8 $\dagger$ $\dagger$
[Uncaptioned image] sketchVLM 70.0 $\pm$ 2.9 68.5 $\pm$ 2.2 92.8 $\pm$ 2.5 75.4 20.4 18.7 11.2 46.69 74.0
[Uncaptioned image] 63.5 $\pm$ 2.5 66.0 $\pm$ 4.7 92.3 $\pm$ 2.3 72.0 19.1 22.4 15.4 $\dagger$ $\dagger$
$\bm{+} +\bm{+} [Uncaptioned image]$ 63.0 62.6 93.3 $\pm$ 3.0 91.7 $\dagger$ $\dagger$ $\dagger$ $\dagger$ 39.0
[Uncaptioned image] 37.0 35.9 50.8 $\pm$ 1.5 48.6 $\dagger$ 198.74 9.0
[Uncaptioned image] 27.0 30.3 62.5 $\pm$ 2.1 68.1 $\dagger$ $\dagger$ $\dagger$ $\dagger$ 0.0
$\dagger$ Model is unable to output this format.
5.1 Grid prompting improves [Uncaptioned image] sketchVLM annotation precision, but hurts [Uncaptioned image] sketchVLM
To understand how each component of our framework affects the model, we run ablations to see how sketching and grid prompting affect [Uncaptioned image] sketchVLM and [Uncaptioned image] sketchVLM .
Experiment   We evaluate four input configurations in single-turn mode: the base image alone, image with grid, image with sketching prompt, and image with both. We report accuracy on VPCT, Ball Drop, and Maze Navigation, and RMSE on Connect-the-Dots (Tab.˜3).
Results   [Uncaptioned image] performs best with both the grid and sketching prompt, with the grid providing a consistent boost to spatial precision (Fig.˜D12), consistent with prior findings [20]. [Uncaptioned image] performs better without the grid, and adding it notably degrades localization on Connect-the-Dots (RMSE increases from 5.92 to 99.34). We therefore report all SketchVLM results in Tab.˜2 with the input grid for [Uncaptioned image] but without the grid for [Uncaptioned image] (Fig.˜D14).
Table 3: Ablation across inputs in single-turn mode. “Sketch” adds strokes/system prompt; “Grid” additionally overlays the coordinate grid. RMSE is reported for Connect-the-Dots while accuracy is reported for the other tasks. [Uncaptioned image] sketchVLM works best with the grid while [Uncaptioned image] sketchVLM works best without the grid
Model Input VPCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] Image 63.5 66.0 92.3 N/A
+ Grid 67.0 67.2 98.0 N/A
+ Sketch 59.0 63.1 88.7 134.40
+ Sketch + Grid + Multi-turn 68.0 63.1 82.1 50.68
+ Sketch + Grid (i.e., [Uncaptioned image] sketchVLM) 70.0 68.5 92.8 46.69
[Uncaptioned image] Image 89.3 83.8 99.3 N/A
+ Grid 90.0 81.3 99.5 N/A
+ Sketch (i.e., [Uncaptioned image] sketchVLM ) 96.0 79.7 98.0 5.92
+ Sketch + Multi-turn 80.0 79.8 98.2 12.37
+ Sketch + Grid 91.0 82.3 99.5 99.34
5.2 SketchVLMs can localize points and connect them in order [Uncaptioned image]
Connecting the dots in an image tests both spatial grounding and the ability to produce multiple coherent strokes in a row.
Experiment   We evaluate the root mean squared error (RMSE) in pixels of the difference between the ground truth position of each of the points and the SketchVLMs’ closest generated point. To evaluate whether models connect points in the correct order, we compare each predicted segment i against all ground-truth segment pairs using MSE. If segment i has lower MSE to a different ground-truth pair than its expected pair (i, i+1), it is counted as an ordering error. Because [Uncaptioned image] and [Uncaptioned image] only produce an image response with no x-y coordinates, we manually evaluate their outputs.
Results   [Uncaptioned image] sketchVLM can accurately output the correct location of up to 35 points with a very low RMSE of only 5.92 (Tab.˜2). [Uncaptioned image] sketchVLM has a higher RMSE of 46.69, but still performs much more reliably than [Uncaptioned image]’s 198.74 RMSE. Regarding ordering, [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image] frequently produce ordering errors as the number of strokes increases, often connecting points out of order and inadvertently altering the input image (Sec.˜C.2). In contrast, [Uncaptioned image] sketchVLM and [Uncaptioned image] sketchVLM correctly order points 74% and 99% of the time, respectively, demonstrating that SketchVLMs can reliably scale to many strokes while maintaining spatial accuracy and logical coherence (Fig.˜D13).
5.3 SketchVLM improves counting accuracy [Uncaptioned image]
InputRefer to caption
Pred: 9 Refer to caption
(a) $\bm{+} +\bm{+} Refer to caption$
Answer: 8 Refer to caption
(b) Refer to caption sketchVLM
Answer: 9 Refer to caption
(c) Refer to caption
Answer: 3 Refer to caption
Figure 5: (a) $\bm{+} +\bm{+} Refer to caption$ generates a different image and predicts an incorrect count. (c) Refer to caption directly outputs only a number without annotations and severely undercounts. In contrast, our Refer to caption sketchVLM (b) outputs the correct answer and produces visual annotations to explain its answer.
Existing VLMs can output point coordinates to mark counted objects, but these points are unlabeled and can be tedious to verify. In contrast, SketchVLMs explicitly place numeric markers on each object, enabling direct visual verification of the predicted count.
Experiment   We measure the accuracy of SketchVLMs and also how well they ground their markers on counting datasets (Sec.˜4.1). We consider a marker correct if it lies within the corresponding ground-truth bounding box, obtained via SAM-3 [7], allowing at most one marker per object.
Results   [Uncaptioned image] sketchVLM exhibits strong consistency between counting accuracy (94.5) and numeric-marker location accuracy (95.9), whereas [Uncaptioned image] sketchVLM achieves high counting accuracy (75.4) but substantially lower numeric-marker location accuracy (51.0), indicating that [Uncaptioned image] sketchVLM often places markers incorrectly despite producing the correct count (Tabs.˜2 and C3). [Uncaptioned image] attains low counting accuracy (48.6) and numeric-marker location accuracy (59.9), indicating limited performance in both aspects (Tabs.˜2 and C3). SketchVLM improves counting accuracy, yielding gains of (+1.5) points for [Uncaptioned image] and (+3.4) points for [Uncaptioned image] (Tab.˜2). The explicit numeric markers further enable direct visual verification of model outputs (Figs.˜5 and D9).
5.4 SketchVLMs localize objects more accurately with pre-defined shape primitives than with free-form annotations [Uncaptioned image]
(a) Refer to caption sketchVLM (b) Refer to caption sketchVLM (c) Refer to caption (d) Refer to caption
Free-form Ovals Free-form Rectangles Oval Primitive
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: When prompted to outline the classes “person” and “sports-ball”, (c) Refer to caption replaces the original image with a newly generated one, whereas SketchVLM in (a) and (b) preserves the original image and draws shapes that accurately align with object boundaries and locations as compared to default (d) Refer to caption.
A design choice we face is whether to have SketchVLMs generate all of their drawings through free-form strokes, or whether to allow them to output shape information (such as the xxyy position of the center of the circle and the length of the radius) and then delegate the actual rendering of the drawing to the SVG conversion. Free-form strokes are more flexible but require point-by-point drawing and demand higher spatial precision. In contrast, pre-defined primitives for rectangles, ovals, etc. can be specified by a few parameters and can be rendered automatically. We compare both approaches in object localization in order to understand how it affects the output of annotations.
Experiment   We compare SketchVLMs’ stroke-based annotations against the baseline models that directly output parameters for shape locations and shape size (Sec.˜4.1). To evaluate oval annotations, we convert the rendered shape into its tight enclosing bounding box before computing metrics. Performance is measured using Average Precision (AP) at an IoU threshold of 0.5.
Results   [Uncaptioned image] sketchVLM with stroke-based rectangles is effective for medium and large objects, but remains limited for small-object detection. Sketch-based rectangles slightly improve performance on medium (+0.6) and large objects (+1.4), but significantly degrade small-object detection (-10.1), reducing overall AP50 from 63.1 to 58.8 (Figs.˜6, D10 and C4).
We observe that stroke-based outputs in SketchVLMs underperform the original model on small objects (Tab.˜C4). We examine detection statistics and prompt ablations and find that SketchVLMs match the original model in precision but exhibit lower recall (Tab.˜C5) and do not significantly change with variations in the drawing prompt (Tab.˜C6). Therefore, in order to get the accuracy of pre-defined shapes as well as the expressiveness of free-form strokes, we allow the model to output both simultaneously, such as in Fig.˜1.
5.5 SketchVLM improves localization accuracy for [Uncaptioned image] but not for [Uncaptioned image] when labeling parts of an object [Uncaptioned image]
Input
(a) Refer to caption
(b) Refer to caption sketchVLM
(c) Refer to caption
Figure 7: Qualitative comparison on the part labeling task. (b) Refer to caption SketchVLM places each part label directly on its corresponding region while preserving the original image, producing more interpretable part annotations than (a) Refer to caption or (c) Refer to caption.
A useful feature of SketchVLMs is pointing at parts of an image and explaining them, for example, labeling engine components in a car maintenance guide (Fig.˜1) or annotating regions in a screenshot (Fig.˜A1). We test SketchVLMs’ ability to produce text that is both semantically correct and spatially well-placed, a capability that underpins step-by-step instructions, explainable visual descriptions, and annotation tools [56, 28, 17].
Experiment   In SketchVLM, the framework automatically renders the predicted text at the predicted location (with adaptive size/color for visibility), whereas the original VLM model outputs only the label and coordinates. For both SketchVLMs and the original VLMs, we prompt the models with a predefined set of valid part labels and require all predictions to be selected strictly from this set. For the original VLMs, we use a Python post-processing pipeline to overlay the predicted label onto the image at the predicted coordinates, where text size and color are manually chosen for consistent visibility across examples. We follow [11] and define boundary dilation with radius $r$ as expanding the ground-truth boundary by $r$ pixels in all directions to allow tolerance in spatial matching.
Results   SketchVLM improves part labeling for [Uncaptioned image] (+1.3) but slightly underperforms [Uncaptioned image] at strict boundary matching (-3.8) (Tab.˜4). For [Uncaptioned image], SketchVLM becomes increasingly robust under boundary dilation [11], achieving higher accuracy as tolerance increases, with remaining errors dominated by wrong-position mistakes (>79%; Tab.˜C7). For [Uncaptioned image], the gap narrows from (-3.8) at $r=0$ to (-0.6) at $r=7$, reaching near parity under modest tolerance (Tab.˜4). These errors largely correspond to small boundary offsets that are visually negligible (Figs.˜7, 8 and D11), where the original models produce more missing-label errors and SketchVLMs more position errors (Tab.˜C7). These findings indicate that SketchVLMs primarily introduce minor spatial shifts rather than semantic labeling failures, and remain competitive under reasonable boundary tolerance.
Table 4: Labels placed by SketchVLMs land very close to the correct location. [Uncaptioned image] sketchVLM is more accurate than the baseline at every tolerance level, and [Uncaptioned image] sketchVLM matches it within a few pixels. Dilation $r$ (px) [Uncaptioned image] [Uncaptioned image] sketchVLM [Uncaptioned image] [Uncaptioned image] sketchVLM 0 64.1 60.3 (-3.8) 19.1 20.4 (+1.3) 3 69.0 66.5 (-2.5) 21.9 23.3 (+1.4) 5 70.8 69.3 (-1.5) 23.9 25.3 (+1.4) 7 71.8 71.2 (-0.6) 25.6 27.2 (+1.6) 10 73.1 73.5 (+0.4) 28.6 30.2 (+1.6) 15 75.1 78.3 (+3.2) 33.7 35.9 (+2.2)    [Uncaptioned image] Figure 8: From $r=0$ (red) to modest dilation ($r=7$, yellow), small boundary offsets become visually negligible.
5.6 SketchVLMs outperform models fine-tuned directly on path-tracing tasks [Uncaptioned image]
(a) Input
(b) Refer to caption sketchVLM
(c) Input
(d) Refer to caption sketchVLM
Input Path: Up, Right, Up, Right, Down
: “… the path is valid.” Refer to caption
Input Path: Up, Up, Right, Down, Down
: “… the path is invalid.” Refer to caption
Figure 9: Models are presented with a blank maze such as (a) or (c) and are asked to verify whether a proposed path from the green square to the red square is feasible through annotations (b) and (d). SketchVLMs correctly verify both valid and invalid paths by drawing the trajectory and marking where an invalid move occurs.
Maze Navigation evaluates spatial reasoning by requiring models to follow a sequence of directions to reach a goal while avoiding obstacles. Given that [Uncaptioned image] and [Uncaptioned image] are trained on a similar maze task, we expect them to perform well.
Experiment   Given a set of directions, the model must sketch out the path while also determining if the path reaches the goal without crossing any walls.
Results   Surprisingly, we find that [Uncaptioned image] has an accuracy of 50.8% and [Uncaptioned image] has an accuracy of 62.5%, both of which are near the random choice baseline of 50%. [Uncaptioned image] + [Uncaptioned image] performs much stronger with an accuracy of 93.3%; however, it sometimes alters the entire image and has other unusual outputs (Fig.˜D4). Both [Uncaptioned image] sketchVLM and [Uncaptioned image] sketchVLM perform well with accuracies of 98.0% and 92.8% respectively. The difference in performance between the default model baseline and the SketchVLM mode is minor (Tab.˜2), but the annotations benefit user verification.
5.7 Fine-tuned sketching models fail to generalize to unseen physics understanding tasks [Uncaptioned image]
Input
Refer to caption sketchVLM
Refer to caption ThinkMorph
Refer to caption ViLaSR
Figure 10: SketchVLM generates the most accurate Ball Drop images compared to other baselines.
In addition to spatial reasoning, we evaluate whether SketchVLM can predict trajectories involving physical dynamics, such as a ball falling and rolling.
Experiment   Given an image with a ball and platforms, the model must sketch the ball’s trajectory and output the container number it lands in.
Results   We find that [Uncaptioned image] and [Uncaptioned image] perform poorly on VPCT with accuracies of 37% and 27%, which are near the random choice baseline of 33.3%. On our Ball Drop dataset, they have accuracies of 35.9% and 30.3%, only slightly above the random choice baseline of 25.0%, showing how these models fail to generalize to tasks that they are not trained on. $\bm{+} +\bm{+} [Uncaptioned image]$ often removes ledges in the image (Fig.˜D1) and performs significantly worse than the baseline [Uncaptioned image] accuracy (Tab.˜2). In contrast, SketchVLMs output coherent annotations and have high accuracy, and [Uncaptioned image] sketchVLM reaches 96.0% accuracy on VPCT and 79.7% on Ball Drop (Fig.˜10).
5.8 SketchVLM has higher annotation–text alignment than image-editing and fine-tuned models
Table 5: VLM-judged annotation–text alignment (higher is better) and annotation quality (1–5, higher is better). Models fine-tuned to generate annotations show lower alignment, while [Uncaptioned image] sketchVLM has the highest alignment and annotation quality.
Model Annotation–text Alignment Annotation Quality (1-5)
VPCT Ball Drop Maze Navigation Mean VPCT Ball Drop Maze Navigation Mean
[Uncaptioned image] sketchVLM 99.0 99.0 88.4 95.5 1.83 1.74 3.20 2.26
[Uncaptioned image] sketchVLM 100.0 98.5 84.2 94.2 3.12 4.28 3.69 3.70
$\bm{+} +\bm{+} [Uncaptioned image]$ 93.0 41.4 80.3 71.6 1.56 2.56 3.68 2.60
[Uncaptioned image] 32.0 45.5 8.2 28.6 1.36 1.28 2.78 1.81
[Uncaptioned image] 54.0 38.4 48.1 46.8 1.62 2.11 1.17 1.63
Experiment   To measure annotation–text alignment, we show each model’s annotations for VPCT, Ball Drop and Maze Navigation (without its text answer) to a VLM judge (Gemini-3-Flash-Preview [12]) and ask it to infer what the final answer should be from the annotations alone.
Results   Despite being training-free, SketchVLMs achieve substantially higher annotation–text alignment than all baselines (Tab.˜5). [Uncaptioned image] sketchVLM and [Uncaptioned image] sketchVLM reach mean alignment scores of 95.5% and 94.2%, respectively, compared to 71.6% for $\bm{+} +\bm{+} [Uncaptioned image]$, 46.8% for [Uncaptioned image], and 28.6% for [Uncaptioned image]. The low alignment of the fine-tuned models means their annotations frequently contradict their own text answers, making them unreliable as visual explanations. In contrast, SketchVLMs’ high alignment means that users can reliably look at the annotations to verify whether the model’s reasoning makes sense, which is the framework’s primary goal.
Input
Prompt: “Draw the path of the ball.”
ViLaSR
Annotation Quality Score: 1
Refer to caption “The physics are completely unrealistic.”
Refer to caption sketchVLM
Annotation Quality Score: 5
Refer to caption “The path is logical, follows gravity, and does not clip through any walls.”
Input
Prompt: “Draw the proposed path: Right, Right, Down, Down.”
ThinkMorph
Annotation Quality Score: 1
Refer to caption “The drawn path completely contradicts the given text path…clips through two solid black walls.”
Refer to caption sketchVLM
Annotation Quality Score: 5
Refer to caption “The sketch follows the proposed path…avoids all solid black walls…There are no logical errors.”
Figure 11: Low-quality annotations from Refer to caption ThinkMorph and Refer to caption ViLaSR may still lead to the correct final answer, but contain logical errors that are harder for users to verify than the high-quality annotations from SketchVLMs.
5.9 [Uncaptioned image] sketchVLM has significantly higher quality annotations than image-editing and fine-tuned models
Beyond faithfulness, an annotation must also be visually clear and logically coherent to be useful. A trajectory that clips through walls, a label that floats in empty space, or uninformative, overlapping annotations all undermine the user’s ability to interpret the model’s reasoning, even if the final text answer is correct (Fig.˜11).
Experiment   As described in Sec.˜4.4, we measure drawing quality using a VLM-as-a-Judge approach with Gemini-3-Flash-Preview [12]. Each annotation is scored on a 1–5 rubric (Appendix˜E) in which the judge is asked to evaluate annotation plausibility and visual clarity.
Results   Based on VLM-as-a-Judge ratings, [Uncaptioned image] sketchVLM achieves the highest mean annotation quality score of 3.70, followed by $\bm{+} +\bm{+} [Uncaptioned image]$ at 2.60, [Uncaptioned image] sketchVLM at 2.26, [Uncaptioned image] at 1.81, and [Uncaptioned image] at 1.63 (Tab.˜5). Similarly, based on human ratings, [Uncaptioned image] sketchVLM achieves the highest mean quality score of 4.14, followed by [Uncaptioned image] sketchVLM at 3.70, Nano Banana Pro at 3.08, [Uncaptioned image] at 1.74, and [Uncaptioned image] at 1.24 (Tab.˜E4). Notably, $\bm{+} +\bm{+} [Uncaptioned image]$ produces annotations that appear visually polished but sometimes contain subtle logical errors (e.g., trajectories passing through solid platforms) that even the VLM judge can miss (Sec.˜E.1). The fine-tuned models [Uncaptioned image] and [Uncaptioned image] score lowest, and manual inspection confirms that their annotations are often incoherent and difficult to interpret (Sec.˜E.1). These low quality scores also correlate with their low task accuracy in Tab.˜2, suggesting that models unable to produce clear annotations also struggle to reason about the underlying tasks.
5.10 VLM judge ratings positively correlate with human judgments
We validate our annotation quality ratings from our VLM judge (Gemini-3-Flash-Preview [12] (Tab.˜F1)) against three human annotators on 2,250 annotations across Ball Drop, VPCT, and Maze Navigation using a quadratic Kappa score and Pearson Correlation. Agreement between humans and the VLM judge is moderate (quadratic Kappa of $0.51\pm 0.02$, Pearson correlation of $0.52\pm 0.01$) compared to high inter-human agreement (quadratic Kappa of $0.84\pm 0.04$, Pearson correlation of $0.85\pm 0.04$). Further details are in Appendix˜E. While the VLM judge is not as consistent as the human evaluators, its ratings are positively correlated with human judgments across all tasks and models. We therefore treat the VLM judge scores as a useful proxy metric for quick and cost-effective evaluation that should be taken into consideration with both task accuracy and annotation–text alignment.
5.11 Single-turn is as accurate as multi-turn but requires significantly fewer turns
Figure 12: Multi-turn example of SketchVLM guiding a user through how to remove an image’s background. At each turn, the model receives a screenshot then annotates the screenshot with labeled arrows and highlights UI elements to indicate the next step.
We want SketchVLMs to be able to visually guide users through tasks that require multiple turns, such as removing the background of a photo (Fig.˜12) or setting up an EC2 instance on AWS (Fig.˜A1). We therefore evaluate how multi-turn generation affects model accuracy compared to single-turn, and identify the most effective configuration for multi-turn interaction.
Experiment   We compare single-turn and multi-turn generation on VPCT, [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. In multi-turn, the model receives all previously drawn strokes rendered onto the input image at each turn. We also test whether providing the text representation of rendered strokes across turns is necessary for maintaining annotation quality.
Results   Single-turn achieves comparable or higher accuracy than multi-turn across all tasks (Tab.˜3), while requiring about 5.92x fewer turns (Tab.˜C9). This demonstrates that SketchVLMs can produce high-quality annotations and accurate answers in a single pass. When we remove the text representation of prior strokes to force the model to rely solely on the rendered image, we observe notable degradation in annotation quality (Fig.˜D15). We therefore report all multi-turn results with both the rendered image and the text history provided to the model.
6 Conclusion
We present SketchVLM, a training-free framework that prompts frontier VLMs to produce editable, non-destructive SVG annotations grounded on the input image. These visual explanations let users verify model reasoning at a glance, something that text-only responses and image-editing baselines fail to reliably provide. Our results show that this approach outperforms sketching models by +28.5 percentage points in accuracy (Tab.˜C1) and +48.3% in annotation quality (Tab.˜C2), while generalizing well to real-world tasks (Appendix˜A).
Limitations and Future Work   SketchVLM works best with [Uncaptioned image] and [Uncaptioned image] and can transfer to strong open-source VLMs like Kimi K2.5 [44] (Tabs.˜C8 and D), but does not perform well on small VLMs that struggle with instruction following like Qwen2.5-VL-7B [3]. Additionally, enabling models to undo and edit strokes could be explored as a way to improve multi-turn performance.
7 Author contribution statement
Brandon Collins (BC), Logan Bolton (LB), and Hung Nguyen (HN) are major contributors who (a) created or curated benchmark datasets and (b) ran experiments.
  • HN curated the Counting Objects, Drawing Shapes around Objects, and Part Labeling benchmarks from existing datasets, led their evaluation, and led the human-versus-VLM agreement study.
  • BC created the Connect-the-Dots benchmark and led method development, inference and evaluation code, ablations, and evaluation on Connect-the-Dots and VPCT.
  • LB created the Ball Drop and Maze Navigation benchmarks, led their evaluation, and ran the open-source model experiments as well as the annotation-quality and annotation-text alignment evaluations.
BC and LB led the writing of the manuscript while all authors contributed to editing and reviewing. BC, LB, and Anh Nguyen (AN) developed the demo. BC led development of the project website, with additional contributions from LB. BC and LB are technical team leads. AN supervised the project.
Acknowledgement
We thank Pooyan Rahmanzadehgervi at Auburn University for feedback and discussions of results. AN was supported by the NSF Grant No. 2145767, and donations from NaphCare Foundation & Adobe Research. LB was supported by the Auburn University URF program. HN was supported by the Auburn University AU PGRF program.
Appendix A Real World Applications
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure A1: Beginning at the top left image, our SketchVLM framework visually guides the user how to set up a free EC2 instance through the notoriously non-user-friendly Amazon Web Services web interface.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure A2: SketchVLM responds via visual annotations while ChatGPT responds with text only.
How do I turn off my hot water?
I’ve got two sticks of RAM. Where do they go?
Where is the ground nutmeg?
Figure A3: SketchVLM can be used in a variety of real-world use cases.
Figure A4: Multi-turn session where the model explains how to set up PyTorch in a Docker container.
Figure A5: Multi-turn session where the model explains how to set up a CI/CD pipeline in GitHub.
Appendix B Dataset Creation
B.1 Connect-the-Dots
We collect 100 connect-the-dots images spanning three subsets with varying dot counts and background clutter.
  1. 1.
    Random dots: We use a Python script with the Pillow image library to generate 3 images each with 4, 5…10 dots randomly placed for a total of 21 images. Each dot has a number that corresponds to the order it should be connected.
  2. 2.
    Outlines: We convert 30 silhouette SVGs from Openclipart [32] into connect-the-dots puzzles. Each SVG was flattened to polylines, simplified with the Douglas–Peucker algorithm [16], and normalized to a unit square. The main contour was then resampled to place 30 evenly spaced dots along the boundary, with number labels slightly offset. We manually filtered out cases with distorted or self-intersecting shapes, which often occurred for concave outlines.
  3. 3.
    Worksheets: We gather 49 images of preexisting connect-the-dot worksheets from online sources. These images differ from the outlined images in how they often contain irrelevant information such as lines at the top of the image for students to mark their name. To obtain the ground truth strokes for these images, we manually annotate the coordinates of each dot.
B.2 Counting
The counting dataset combines three sources: CountBench [5], TallyQA [1], and Pixmo-Count [14]. The CountBench and TallyQA subsets together contain 746 samples, covering object counts ranging from 0 to 10. Pixmo-Count contributes 443 samples after removing unsuitable cases from the original 526-image test split and similarly includes object counts from 1 to 10.
B.3 Drawing Shapes around Objects
The testing dataset consists of 1,000 carefully selected images from the 5,000 COCO validation images, chosen to ensure a balanced distribution of object counts across classes and object sizes (small, medium, and large).
B.4 Part Labeling
We carefully selected images from two datasets, PACO [37] and Pascal-Part [10]. The selected images satisfy the following criteria:
  1. 1.
    Each image contains only one object corresponding to the target class name.
  2. 2.
    The object’s size occupies at least 10% of the total image area.
  3. 3.
    Each selected object has at least four part labels annotated.
  4. 4.
    The dataset maintains a balanced distribution of objects across different classes.
After selection, the final dataset used for the part labeling task consists of 985 images covering 52 class names.
B.5 Maze Navigation
Given a start point, an end point, and a set of direction commands (e.g., Up, Down, Left, Right), the model must determine if the path reaches the goal without crossing any border walls. We create 200 unique 3x3 grids where the shortest path length between the starting green square and ending red square varies from 3 to 8 steps. For each maze, we take the ground truth path and randomly change one of the direction steps in order to make an invalid path (for example, Left, Right, Down could be changed to Right, Right, Down) (Fig.˜9).
B.6 Ball Drop
We evaluate our framework using Visual Physics Comprehension Test (VPCT [8]), which consists of 100 hand-crafted images where the model must determine which of the buckets the ball will fall into after it is dropped. While VPCT does provide a simple way to evaluate physics understanding of VLMs, it does not contain any ground truth data of the trajectory of the ball. Therefore, in order to evaluate how well SketchVLM models can draw the true trajectory of the ball paths, we generate our own benchmark (Ball Drop). We simulate the trajectory of the ball using PHYRE [4] to obtain ground truth ball trajectory data. In contrast to VPCT, our Ball Drop benchmark is synthetically generated. We generate 198 unique images, with an equal number of images containing 1, 2, and 3 randomly placed lines. We randomize the ball’s X position and fix its Y position near the top. There are four containers at the bottom of the image compared to VPCT’s three containers, making it harder to guess answers correctly.
Appendix C Additional Task Results
C.1 Combined Results
Table C1: SketchVLM improves visual reasoning task accuracy by +28.5 points over alternative sketching approaches. Averages are computed over VPCT, Ball Drop, Maze, and Counting tasks.
Category Model VPCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Avg $\Delta$
SketchVLM [Uncaptioned image] sketchVLM 96.0 79.7 98.0 94.5 84.4 +28.5
[Uncaptioned image] sketchVLM 70.0 68.5 92.8 75.4
Other Sketching $\bm{+} +\bm{+} [Uncaptioned image]$ 63.0 62.6 93.3 91.7 55.9
[Uncaptioned image] 37.0 35.9 50.8 48.6
[Uncaptioned image] 27.0 30.3 62.5 68.1
Table C2: SketchVLM produces higher quality annotations, scoring +48.3% above alternative sketching approaches on a 1–5 VLM-judged drawing quality scale.
Category Model VPCT [Uncaptioned image] [Uncaptioned image] Avg $\Delta$
SketchVLM (Ours) [Uncaptioned image] sketchVLM 3.12 4.28 3.69 2.98 +48.3%
[Uncaptioned image] sketchVLM 1.83 1.74 3.20
Other Sketching $\bm{+} +\bm{+} [Uncaptioned image]$ 1.56 2.56 3.68 2.01
[Uncaptioned image] 1.36 1.28 2.78
[Uncaptioned image] 1.62 2.11 1.17
C.2 Connect-the-Dots
Figure C1: Gemini-2.5-Flash-Image frequently adds non-existent points to the image or does not correctly follow the proper order of the connected dots. Nano Banana Pro is significantly better than Gemini-2.5-Flash-Image, but still only completes the task without any errors 37% of the time.
Random Dots (n=21) Worksheets (n=49) Outlines (n=30) Total (n=100)
Model No Grid With Grid $\Delta$ No Grid With Grid $\Delta$ No Grid With Grid $\Delta$ No Grid With Grid $\Delta$
Gemini-2.5-Pro 1212.54 429.86 -64.55% 14769.86 1188.59 -91.95% 7848.27 1597.67 -79.65% 9846.22 1151.49 -88.31%
Gemini-3-Pro 15.70 584.32 +3621.78% 48.60 1256.37 +2485.12% 25.00 30436.03 +121644.11% 34.61 9869.14 +28414.46%
GPT-5 (low) 11988.22 402.50 -96.64% 16816.74 4627.17 -72.48% 24355.01 2885.04 -88.15% 18063.82 2179.59 -87.93%
GPT-5 (med) 3098.15 427.02 -86.22% 15060.98 4627.17 -69.27% 22520.39 2885.04 -87.19% 14785.98 3222.40 -78.21%
GPT-5 (high) 2415.63 419.35 -82.64% 14859.81 2219.47 -85.06% 20789.13 3362.99 -83.82% 14024.76 2183.90 -84.43%
Qwen3-8B 12667.46 17728.22 +40.00% 80952.85 157661.81 +94.76% 484212.61 542690.99 +12.08% 187590.15 243784.56 +29.96%
Qwen3-235B 2857.16 845.90 -70.38% 21472.98 26530.78 +23.56% 34420.65 227671.27 +561.40% 21447.25 81479.11 +280.00%
ViLaSR 26658.05 N/A N/A 37153.54 N/A N/A 52312.84 N/A N/A 39497.28 N/A N/A
Kimi K2.5 947.41 386.91 -59.16% 8874.52 3256.55 -63.30% 31596.88 6010.74 -80.98% 14026.53 3480.18 -75.19%
Figure C2: Connect-the-Dots Mean MSE with categories as columns. Each entry shows the mean MSE for with grid and without grid, and the percent change $\Delta$ (negative is better).
C.3 Counting
Table C3: [Uncaptioned image] sketchVLM achieves high text-location accuracy, indicating that predicted counts are well aligned with target objects, whereas [Uncaptioned image] sketchVLM shows substantially weaker grounding.
Model Accuracy (%)
[Uncaptioned image] sketchVLM 95.9
[Uncaptioned image] sketchVLM 51.0
[Uncaptioned image] 59.9
C.4 Drawing Shapes around Objects
Table C4: Sketch-based localization improves accuracy for medium and large objects, while reducing performance on small objects, compared to coordinate-based bounding boxes (AP50).
[Uncaptioned image] [Uncaptioned image] sketchVLM [Uncaptioned image] sketchVLM
Output format Bounding box Rectangle Oval
AP50 (all) 63.1 58.8 (-4.3) 55.4 (-7.7)
AP50 (small) 35.6 25.5 (-10.1) 26.4 (-9.2)
AP50 (medium) 62.6 63.2 (+0.6) 55.4 (-7.2)
AP50 (large) 77.7 79.1 (+1.4) 77.6 (-0.1)
Table C5: [Uncaptioned image] and [Uncaptioned image] SketchVLM achieve the same precision; however, SketchVLM has lower recall than the baseline because it produces fewer true positive (TP) detections.
GT [Uncaptioned image] [Uncaptioned image] SketchVLM TP FP FN
P R AP P R AP [Uncaptioned image] [Uncaptioned image] sketchVLM [Uncaptioned image] [Uncaptioned image] sketchVLM [Uncaptioned image] [Uncaptioned image] sketchVLM
1060 18.7 47.9 33.1 18.7 35.4 26.0 508 375 2204 1629 552 685
Table C6: Different prompt variations do not improve AP performance for the drawing shape task.
[Uncaptioned image] [Uncaptioned image] SketchVLM [Uncaptioned image] SketchVLM [Uncaptioned image] SketchVLM [Uncaptioned image] SketchVLM
Draw rectangles Draw all visible/part visible Scan region before draw Count object before draw
Output shape Bounding box Rectangles Rectangles Rectangles Rectangles
AP50 (all) 63.1 58.8 57.1 59.0 56.2
AP50 (small) 35.6 25.5 27.5 26.8 23.2
AP50 (medium) 62.6 63.2 60.7 59.9 58.9
AP50 (large) 77.7 79.1 77.8 80.7 78.3
C.5 Part Labeling
Table C7: Error type breakdown for labeling and part labeling with [Uncaptioned image] and [Uncaptioned image] original models and SketchVLM.
Wrong Type [Uncaptioned image] (%) [Uncaptioned image] SketchVLM(%) [Uncaptioned image] (%) [Uncaptioned image] SketchVLM(%)
Missing Label 10.8 3.6 0.69 0.32
Wrong Position 25.1 36.1 80.17 79.32
Total 35.9 39.7 80.86 79.64
C.6 Additional Model Results
Table C8: Ablation across inputs in single-turn mode for additional models. “Sketch” adds strokes/system prompt; “Grid” additionally overlays the coordinate grid. RMSE is reported for Connect-the-Dots while accuracy is reported for the other tasks. “Order Accuracy” is the percentage of connect-the-dots samples with correct point ordering (higher is better).
Model Input VPCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Order Accuracy
GPT-5 (med) Image 63.3 72.2 121.60
+ Sketch + Grid 77.0 71.7 56.76
Gemini-2.5-Flash Image 38.0 31.1 57.5
+ Sketch + Grid 51.0 50.7 70.0
Gemini-2.5-Pro Image 50.0 43.5 72.2 99.23
+ Sketch + Grid 61.0 57.5 81.5 33.93
Qwen-2.5VL-7B Image 37.0 22.3 46.3
+ Sketch + Grid 34.0 1.0 34.0
Kimi K2.5 Image 44.0 50.50 93.5 118.43 59%
+ Grid 55.0 58.08 97.0
+ Sketch 56.0 48.48 88.00
+ Sketch + Grid 56.0 50.50 89.75 58.99 67%
Table C9: Average number of turns per task group in multi-turn evaluation.
Task Gemini-3-Pro GPT-5 (low)
VPCT 5.10 6.18
[Uncaptioned image] 5.00 2.58
[Uncaptioned image] 16.61 19.00
[Uncaptioned image] 3.42 4.29
Overall 5.91 5.92
Appendix D Qualitative Samples
D.1 Ball Drop
Figure D1: Qualitative examples of models on VPCT. Gemini-3-Pro-Preview in single-turn produces the most accurate annotations, while NanoBanana Pro, ThinkMorph and ViLaSR often draw paths that cross walls and provide the wrong answer.
Figure D2: Qualitative examples of models on our Ball Drop dataset. Gemini-3-Pro-Preview in single-turn produces the most accurate annotations, while NanoBanana Pro, ThinkMorph and ViLaSR often draw paths that cross walls and provide the wrong answer.
Figure D3: The SketchVLM framework boosts VLM accuracy on the Ball Drop task while also letting them draw ball trajectory paths that closely simulate the ground truth data.
D.2 Maze Navigation
Figure D4: Qualitative examples of models on valid paths in Maze Navigation.
Figure D5: Qualitative examples of models on invalid paths in Maze Navigation.
D.3 Connect Dots
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption Kimi
MSE: 1906
Refer to caption Kimi + Grid
MSE: N/A
Refer to caption Qwen3-235B
MSE: 1960
Refer to caption Qwen3-235B + Grid
MSE: 658
Refer to caption Gemini-2.5-Pro
MSE: 3419
Refer to caption Gemini-2.5-Pro + Grid
MSE: 299
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption GPT-5 (low)
MSE: 34784
Refer to caption GPT-5 (low) + Grid
MSE: 387
Refer to caption GPT-5 (med)
MSE: 5799
Refer to caption GPT-5 (med) + Grid
MSE: 617
Refer to caption GPT-5 (high)
MSE: 2545
Refer to caption GPT-5 (high) + Grid
MSE: 390
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption ViLaSR
MSE: 5951
Refer to caption ThinkMorph
MSE: N/A
Refer to caption Gemini-3-Pro
MSE: 3
Refer to caption Gemini-3-Pro + Grid
MSE: 537
Refer to caption Gemini-3-Pro (multi)
MSE: 5
Refer to caption GPT-5 (low) (multi)
MSE: 467
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption Kimi
MSE: 1279
Refer to caption Kimi + Grid
MSE: 479
Refer to caption Qwen3-235B
MSE: 417
Refer to caption Qwen3-235B + Grid
MSE: 791
Refer to caption Gemini-2.5-Pro
MSE: 1414
Refer to caption Gemini-2.5-Pro + Grid
MSE: 542
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption GPT-5 (low)
MSE: 1647
Refer to caption GPT-5 (low) + Grid
MSE: 428
Refer to caption GPT-5 (med)
MSE: 866
Refer to caption GPT-5 (med) + Grid
MSE: 364
Refer to caption GPT-5 (high)
MSE: 757
Refer to caption GPT-5 (high) + Grid
MSE: 428
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption ViLaSR
MSE: 275029
Refer to caption ThinkMorph
MSE: N/A
Refer to caption Gemini-3-Pro
MSE: 15
Refer to caption Gemini-3-Pro + Grid
MSE: 601
Refer to caption Gemini-3-Pro (multi)
MSE: 5
Refer to caption GPT-5 (low) (multi)
MSE: 456
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption Kimi
MSE: 573
Refer to caption Kimi + Grid
MSE: 644
Refer to caption Qwen3-235B
MSE: 3420
Refer to caption Qwen3-235B + Grid
MSE: 563
Refer to caption Gemini-2.5-Pro
MSE: 957
Refer to caption Gemini-2.5-Pro + Grid
MSE: 624
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption GPT-5 (low)
MSE: 3137
Refer to caption GPT-5 (low) + Grid
MSE: 412
Refer to caption GPT-5 (med)
MSE: 3369
Refer to caption GPT-5 (med) + Grid
MSE: 393
Refer to caption GPT-5 (high)
MSE: 4506
Refer to caption GPT-5 (high) + Grid
MSE: 528
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption ViLaSR
MSE: N/A
Refer to caption ThinkMorph
MSE: N/A
Refer to caption Gemini-3-Pro
MSE: 11
Refer to caption Gemini-3-Pro + Grid
MSE: 698
Refer to caption Gemini-3-Pro (multi)
MSE: 4
Refer to caption GPT-5 (low) (multi)
MSE: 501
Figure D6: Connect-the-Dots qualitative comparisons on random dots. Each item spans three rows: (top) Kimi/Qwen3-235B/Gemini-2.5-Pro, (middle) GPT-5 (low/med/high), (bottom) multi-turn variants (Gemini-3-Pro and GPT-5 (low)), with ViLaSR and ThinkMorph added to the third row. Each cell shows the overlay and MSE.
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption Kimi
MSE: 3446
Refer to caption Kimi + Grid
MSE: 1294
Refer to caption Qwen3-235B
MSE: 1243
Refer to caption Qwen3-235B + Grid
MSE: 817
Refer to caption Gemini-2.5-Pro
MSE: 2626
Refer to caption Gemini-2.5-Pro + Grid
MSE: 377
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption GPT-5 (low)
MSE: 11618
Refer to caption GPT-5 (low) + Grid
MSE: 878
Refer to caption GPT-5 (med)
MSE: 7432
Refer to caption GPT-5 (med) + Grid
MSE: 1033
Refer to caption GPT-5 (high)
MSE: 5420
Refer to caption GPT-5 (high) + Grid
MSE: 651
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption ViLaSR
MSE: 275029
Refer to caption ThinkMorph
MSE: N/A
Refer to caption Gemini-3-Pro
MSE: 6
Refer to caption Gemini-3-Pro + Grid
MSE: 637
Refer to caption Gemini-3-Pro (multi)
MSE: 173
Refer to caption GPT-5 (low) (multi)
MSE: 959
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption Kimi
MSE: 1595
Refer to caption Kimi + Grid
MSE: 5989
Refer to caption Qwen3-235B
MSE: 3422
Refer to caption Qwen3-235B + Grid
MSE: 36302
Refer to caption Gemini-2.5-Pro
MSE: 4162
Refer to caption Gemini-2.5-Pro + Grid
MSE: 823
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption GPT-5 (low)
MSE: 32273
Refer to caption GPT-5 (low) + Grid
MSE: 1119
Refer to caption GPT-5 (med)
MSE: 26805
Refer to caption GPT-5 (med) + Grid
MSE: 1786
Refer to caption GPT-5 (high)
MSE: 34658
Refer to caption GPT-5 (high) + Grid
MSE: 1521
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption ViLaSR
MSE: 671536
Refer to caption ThinkMorph
MSE: N/A
Refer to caption Gemini-3-Pro
MSE: 56
Refer to caption Gemini-3-Pro + Grid
MSE: 3787
Refer to caption Gemini-3-Pro (multi)
MSE: 104
Refer to caption GPT-5 (low) (multi)
MSE: 1982
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption Kimi
MSE: 1092
Refer to caption Kimi + Grid
MSE: N/A
Refer to caption Qwen3-235B
MSE: 304443
Refer to caption Qwen3-235B + Grid
MSE: 26515
Refer to caption Gemini-2.5-Pro
MSE: 3728
Refer to caption Gemini-2.5-Pro + Grid
MSE: 1824
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption GPT-5 (low)
MSE: 50688
Refer to caption GPT-5 (low) + Grid
MSE: 3109
Refer to caption GPT-5 (med)
MSE: 23417
Refer to caption GPT-5 (med) + Grid
MSE: 3798
Refer to caption GPT-5 (high)
MSE: 47798
Refer to caption GPT-5 (high) + Grid
MSE: 6210
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption ViLaSR
MSE: 1222760
Refer to caption ThinkMorph
MSE: N/A
Refer to caption Gemini-3-Pro
MSE: 25
Refer to caption Gemini-3-Pro + Grid
MSE: 3712
Refer to caption Gemini-3-Pro (multi)
MSE: 32
Refer to caption GPT-5 (low) (multi)
MSE: 5732
Figure D7: Connect-the-Dots (worksheets) qualitative comparisons for connect-the-dot worksheets. Each item spans three rows: (top) Kimi/Qwen3-235B/Gemini-2.5-Pro, (middle) GPT-5 (low/med/high), (bottom) multi-turn variants (Gemini-3-Pro and GPT-5 (low)), with ViLaSR and ThinkMorph added to the third row.
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption Kimi
MSE: 11024
Refer to caption Kimi + Grid
MSE: 1503
  Qwen3-235B
MSE: N/A
Refer to caption Qwen3-235B + Grid
MSE: N/A
Refer to caption Gemini-2.5-Pro
MSE: 2685
Refer to caption Gemini-2.5-Pro + Grid
MSE: 1080
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption GPT-5 (low)
MSE: 22042
Refer to caption GPT-5 (low) + Grid
MSE: 2362
Refer to caption GPT-5 (med)
MSE: 27095
Refer to caption GPT-5 (med) + Grid
MSE: 2473
Refer to caption GPT-5 (high)
MSE: 30159
Refer to caption GPT-5 (high) + Grid
MSE: 1134
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption ViLaSR
MSE: 2372280
Refer to caption ThinkMorph
MSE: N/A
Refer to caption Gemini-3-Pro
MSE: 23
Refer to caption Gemini-3-Pro + Grid
MSE: 1370
Refer to caption Gemini-3-Pro (multi)
MSE: 17
Refer to caption GPT-5 (low) (multi)
MSE: 1737
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption Kimi
MSE: 1290
Refer to caption Kimi + Grid
MSE: 1595
Refer to caption Qwen3-235B
MSE: 19328
Refer to caption Qwen3-235B + Grid
MSE: 34308
Refer to caption Gemini-2.5-Pro
MSE: 6867
Refer to caption Gemini-2.5-Pro + Grid
MSE: 1261
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption GPT-5 (low)
MSE: 24565
Refer to caption GPT-5 (low) + Grid
MSE: 8363
Refer to caption GPT-5 (med)
MSE: 30512
Refer to caption GPT-5 (med) + Grid
MSE: 10865
Refer to caption GPT-5 (high)
MSE: 32064
Refer to caption GPT-5 (high) + Grid
MSE: 2765
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption ViLaSR
MSE: N/A
Refer to caption ThinkMorph
MSE: N/A
Refer to caption Gemini-3-Pro
MSE: 20
Refer to caption Gemini-3-Pro + Grid
MSE: 1670
Refer to caption Gemini-3-Pro (multi)
MSE: 23
Refer to caption GPT-5 (low) (multi)
MSE: 3047
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption Kimi
MSE: 2984
Refer to caption Kimi + Grid
MSE: 7336
Refer to caption Qwen3-235B
MSE: 686
Refer to caption Qwen3-235B + Grid
MSE: 534622
Refer to caption Gemini-2.5-Pro
MSE: 7418
Refer to caption Gemini-2.5-Pro + Grid
MSE: 664
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption GPT-5 (low)
MSE: 24851
Refer to caption GPT-5 (low) + Grid
MSE: 1840
Refer to caption GPT-5 (med)
MSE: 20791
Refer to caption GPT-5 (med) + Grid
MSE: 1755
Refer to caption GPT-5 (high)
MSE: 5917
Refer to caption GPT-5 (high) + Grid
MSE: 2447
Refer to caption Source Refer to caption GT (MSE: 0) Refer to caption ViLaSR
MSE: 426532
Refer to caption ThinkMorph
MSE: N/A
Refer to caption Gemini-3-Pro
MSE: 15
Refer to caption Gemini-3-Pro + Grid
MSE: 2702
Refer to caption Gemini-3-Pro (multi)
MSE: 28
Refer to caption GPT-5 (low) (multi)
MSE: 2327
Figure D8: Connect-the-Dots (outlines) qualitative comparisons for connect-the-dot outlines. Each item spans three rows: (top) Kimi/Qwen3-235B/Gemini-2.5-Pro pairs, (middle) GPT-5 (low/med/high), (bottom) multi-turn variants (Gemini-3-Pro and GPT-5 (low)), with ViLaSR and ThinkMorph added to the third row.
D.4 Counting
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure D9: Qualitative comparison on counting task between Refer to caption sketchVLM , Refer to caption sketchVLM, Refer to caption and Refer to caption.
D.5 Drawing Shape
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure D10: Qualitative comparison on drawing shape tasks between Refer to caption, Refer to caption sketchVLM , and Refer to caption.
D.6 Part Labeling
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure D11: Qualitative comparison on part labeling tasks between Refer to caption, Refer to caption sketchVLM , and Refer to caption.
D.7 Connect-the-Dots: Grid versus No Grid
Source
Refer to caption Sketch without Grid
Refer to caption Sketch with Grid
Figure D12: Appending a reference coordinate grid to the edge of the input image allows Refer to caption to be more precise, but does not help Refer to caption.
D.8 Connect-the-Dots: Bézier Curves versus Lines
Straight Lines: 12 strokes
Bézier Curves: 2 strokes
Straight Lines: 9 strokes
Bézier Curves: 1 stroke
Straight Lines: 17 strokes
Bézier Curves: 9 strokes
Straight Lines: 9 strokes
Bézier Curves: 3 strokes
Straight Lines: 19 strokes
Bézier Curves: 1 stroke
Figure D13: Using Bézier curves instead of straight lines allows the model to connect the dots in fewer strokes, and it can draw less jagged shapes, leading to a more aesthetically pleasing result.
D.9 Gemini-3-Pro-Preview Coordinate Systems
Figure D14: Comparing Gemini-3-Pro-Preview annotations under two coordinate systems. The 0–2000 system (vs. the 0–1000 system the model is typically used to) often preserves the overall shape but produces annotations that are compressed or shifted. In most cases the model adapts to the new coordinate system, but these failure cases lead to a noticeable decrease in performance. We hypothesize this change is what causes Gemini-3-Pro-Preview to also perform worse with the grid (which has a different coordinate system).
D.10 Multi-turn Ablation
Multiturn: Prompts + Image + Text Annotations sent back each turn
Refer to captionRefer to caption
Multiturn: Prompts + Image sent back each turn
Refer to captionRefer to caption
Figure D15: When testing Gemini-3-Pro-Preview in the multi-turn setting, not sending back the text representation of prior annotations degrades drawing quality: the model often attempts to redraw earlier strokes and fails to properly connect new strokes to the existing trajectory.
Appendix E VLM-Judge Details
We use a VLM judge (Gemini-3-Flash-Preview [12]) to evaluate annotation quality and annotation–text alignment. To verify that this automated metric is meaningful, we measure how well VLM judge ratings agree with human annotators.
Experiment   Three independent human annotators each rate 50 annotations across the Ball Drop, VPCT, and Maze Navigation tasks from all 5 models ([Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]), totaling 2,250 labeled images. Both human and VLM judges use the same 1–5 grading rubric (Sec.˜E.2). We report agreement using quadratically weighted Cohen’s $\kappa$, which penalizes larger disagreements more heavily on ordinal scales [21], and Pearson correlation, following recent VLM-judge benchmarks [24].
Results   Human–human agreement is high, with $\kappa=0.84\pm 0.04$ and Pearson $r=0.85\pm 0.04$ (Tab.˜E1), indicating strong consistency between annotators. Detailed per-dataset agreement statistics are reported in Tab.˜E3.
Human–VLM agreement is moderate, with $\kappa=0.51\pm 0.02$ and Pearson $r=0.52\pm 0.01$, with task-level results shown in Tab.˜E2. The VLM judge’s ratings are positively correlated with human judgments across all tasks and models, making it a useful proxy for large-scale evaluation. However, the gap between human–human and human–VLM agreement indicates that the VLM judge can miss subtle errors, particularly logical violations such as trajectories clipping through walls (Sec.˜E.1).
We therefore use the VLM judge for cost-effective large-scale comparison across models, while acknowledging that human annotation remains more reliable for fine-grained evaluation.
Table E1: Agreement is measured by quadratic $\kappa$ and Pearson correlation (mean $\pm$ std). Human annotators are highly consistent, while the VLM judge shows moderate alignment with human judgment.
Comparison Kappa (Quadratic) Pearson
Human–Human $0.84\pm 0.04$ $0.85\pm 0.04$
Human–VLM Judge $0.51\pm 0.02$ $0.52\pm 0.01$
Table E2: Human–VLM agreement measured by quadratically weighted Cohen’s $\kappa$ and Pearson correlation. The VLM judge achieves consistent moderate agreement with all annotators, with higher reliability on valid trajectories and lower agreement when annotations violate maze constraints.
Annotator 1 Annotator 2 Annotator 3
Dataset $\kappa_{quad}$ Pearson $\kappa_{quad}$ Pearson $\kappa_{quad}$ Pearson
ball-path 0.53 0.53 0.51 0.51 0.51 0.52
maze-invalid 0.40 0.41 0.44 0.46 0.42 0.44
maze-valid 0.53 0.58 0.60 0.64 0.63 0.65
vpct 0.42 0.49 0.41 0.44 0.46 0.50
Table E3: Human–human agreement across datasets. Annotators show consistently high agreement ($\kappa\approx 0.85$$0.92$), indicating reliable human evaluation, with slightly lower agreement on maze-invalid cases.
Annotator 1 vs Annotator 2 Annotator 1 vs Annotator 3 Annotator 2 vs Annotator 3
Dataset $\kappa_{quad}$ Pearson $\kappa_{quad}$ Pearson $\kappa_{quad}$ Pearson
ball-path 0.90 0.90 0.90 0.90 0.85 0.85
maze-invalid 0.80 0.82 0.50 0.52 0.53 0.53
maze-valid 0.91 0.92 0.91 0.92 0.92 0.93
vpct 0.89 0.90 0.88 0.89 0.86 0.87
Table E4: Human evaluation score (1–5, higher is better). We report mean $\pm$ standard deviation for each task and the overall mean across all tasks. Best mean in each column is bolded. SketchVLMs have higher mean annotation quality than other annotation models.
Model Human Score (1–5)
VPCT Ball Drop Maze Navigation (Invalid) Maze Navigation (Valid) Mean
[Uncaptioned image] sketchVLM 3.18 $\pm$ 1.27 3.24 $\pm$ 1.08 4.13 $\pm$ 1.24 4.45 $\pm$ 1.22 3.70 $\pm$ 1.32
[Uncaptioned image] sketchVLM 4.56 $\pm$ 0.85 3.79 $\pm$ 1.30 3.92 $\pm$ 1.27 4.36 $\pm$ 1.07 4.14 $\pm$ 1.18
$\bm{+} +\bm{+} [Uncaptioned image]$ 2.44 $\pm$ 1.30 2.94 $\pm$ 1.56 2.77 $\pm$ 1.41 4.29 $\pm$ 1.36 3.08 $\pm$ 1.57
[Uncaptioned image] 1.26 $\pm$ 0.88 1.01 $\pm$ 0.12 3.02 $\pm$ 1.34 1.96 $\pm$ 1.05 1.74 $\pm$ 1.20
[Uncaptioned image] 1.44 $\pm$ 1.01 1.10 $\pm$ 0.38 1.17 $\pm$ 0.45 1.25 $\pm$ 0.79 1.24 $\pm$ 0.72
E.1 Qualitative Examples
Figure E1: Comparison between the original sketching model’s answer and the judge’s answer for valid paths.
Figure E2: Comparison between the original sketching model’s answer and the judge’s answer for invalid paths.
Figure E3: VLM-Judge quality scores for VPCT.
Figure E4: VLM-Judge quality scores for valid paths in the Maze Navigation task.
Figure E5: VLM-Judge quality scores for invalid paths in the Maze Navigation task.
E.2 Rubric Prompts
Rubric for VPCT and Ball Drop Quality
Figure E6: Rubric for VPCT and Ball Drop Quality
Rubric for Path Navigation Quality
Figure E7: Rubric for Maze Navigation Quality
Appendix F Model Settings and Prompts
F.1 API Settings
Table F1: Model inference settings across providers.
Model Details
GPT-5 Model: gpt-5
Temperature: Default
Reasoning Effort: Low, Medium, High
Using OpenAI Responses API
Gemini 2.5 Flash Model: gemini-2.5-flash
Temperature: Default
Using Google AI Studio API
Gemini 2.5 Pro Model: gemini-2.5-pro
Temperature: Default
Using Google AI Studio API
Gemini 3 Pro Model: google/gemini-3-pro-preview
Temperature: Default
Reasoning Effort: Default
Using OpenRouter API (Google AI Studio provider)
Gemini 3 Flash Model: google/gemini-3-flash-preview
Temperature: Default
Reasoning Effort: Default
Using OpenRouter API (Google AI Studio provider)
Nano Banana Pro Model: google/gemini-3-pro-image-preview
Temperature: Default
Output Mode: Image Only
Using OpenRouter API (Google AI Studio provider)
Qwen 2.5 VL 7B Model: qwen/qwen-2.5-vl-7b-instruct
Temperature: Default
Using OpenRouter API (Hyperbolic provider)
ThinkMorph Model: ThinkMorph/ThinkMorph-7B
Temperature: Default
Ran locally using inference code provided at
https://github.com/ThinkMorph/ThinkMorph.
ViLaSR Model: inclusionAI/ViLaSR
Temperature: Default
Ran locally using inference code provided at
https://github.com/AntResearchNLP/ViLaSR/.
F.2 SketchVLM System Prompt
System prompt (1/3)
Figure F1: System Prompt (1/3). Used for all SketchVLM models in single-turn and multi-turn.
System prompt (2/3)
Figure F2: System Prompt (2/3), continued from Fig.˜F1.
System prompt (3/3)
Figure F3: System Prompt (3/3), continued from Fig.˜F2. Parameters are res_x, res_y, and origin. res_x controls how many x coordinates there are. res_y controls how many y coordinates there are. origin controls whether the origin is at the top left or bottom left.
One-stroke guard (multiturn)
[Mode: stepwise]
You are in stepwise mode. On this turn you output EXACTLY ONE stroke block:
<answer>
<strokes>
<sN>...</sN>
</strokes>
</answer>
Do NOT output any other <sM> blocks, no <final_answer>, no explanations.
If the drawing is already complete and no further stroke is needed, output an empty <answer> with NO <strokes> block.
Stop immediately after </answer>.
Final-answer guard (multiturn)
[Mode: stepwise]
All strokes have already been provided. On this turn output ONLY:
<final_answer> ... </final_answer>
Do not output the previous strokes again. Stop immediately after </final_answer>.
Figure F4: During multi-turn, when the model must draw a stroke, we append the "One-stroke guard" to clarify to the model that it should only draw one stroke with no final answer. After the model has decided that no more strokes are needed (it does this by outputting no strokes on a turn), we run one more turn where we prompt the model with "Final-answer guard."
F.3 SketchVLM Output Example
Stroke annotations (XML) <s1> <points>’x500y100’,’x500y270’,’x500y270’,’x350y400’</points> <t_values>0.00,0.50,0.50,1.00</t_values> <id>path_1</id> </s1> <s2> <points>’x350y400’,’x325y470’,’x300y540’</points> <t_values>0.00,0.50,1.00</t_values> <id>drop_1</id> </s2> <s3> <points>’x300y540’,’x390y470’,’x500y670’</points> <t_values>0.00,0.50,1.00</t_values> <id>path_2</id> </s3> <s4> <points>’x500y670’,’x570y690’,’x640y710’</points> <t_values>0.00,0.50,1.00</t_values> <id>path_bounce</id> </s4> <s5> <points>’x640y710’,’x700y850’,’x730y950’</points> <t_values>0.00,0.60,1.00</t_values> <id>drop_2</id> </s5> <final_answer>3</final_answer>
Figure F5: Example sketch output and annotations for prompting SketchVLM with Gemini-3-Pro-Preview on VPCT. Each stroke is colored differently for viewing purposes.
Appendix G Other Baselines
G.1 Baselines
Image-Editing Models.
To benchmark the image-editing model Nano Banana Pro (commonly referred to as Nano Banana Pro), we first prompt Nano Banana Pro to generate a sketch for the input image. We then provide the resulting edited image to Gemini-3-Pro-Preview to produce the final task answer. For Connect-the-Dots, we additionally report a manual evaluation of Nano Banana Pro, since its outputs do not always map cleanly to our structured sketch representation.
Fine-tuned Sketching Models.
For fine-tuned sketching models such as ViLaSR and ThinkMorph, we use the same task prompts as for SketchVLM. Because these models are trained to always output a sketch, they do not have a meaningful “no-sketch” baseline mode; therefore, we omit baseline VQA accuracy for these models in the main results and report only their sketch-conditioned performance.